OpenStudy (anonymous):

You are lying 120 ft away from a tree that is 50 ft tall. You look up at the top of the tree. Approximately how far is your head from the top of the tree?

OpenStudy (amistre64):

the distance of the hypotenuse of the right triangle that is presumably formed by your head being 120 feet from the base of the tree and the height of the tree being 50 ft

OpenStudy (amistre64):

120^2 + 50^ = c^2 14400 + 2500 = c^2 16900 = c^2 sqrt(16900) = c 130 ft